why energy is a scalar quantity?
whereas it flows in a particular direction in magnets and transfer of heat

Grade:11

8 Answers

Akanksha Singh
7 years ago
we can say that energy consumed is equal to the work done. 
energy = work= F*s, here both the force and the time is vector quantity. A
Akanksha Singh
7 years ago
 
we can say that energy consumed is equal to the work done. 
Energy = Work= F*s, here both the force and the time is vector quantity. And we know that the product of two vector quantities give the scalar quantity. So we can say that energy is a scalar quantity.
DILIP
7 years ago
Kinetic energy is a scalar because it is defined to be a scalar.

The reason why scientists define certain words to mean certain things is that some definitions are more useful than others. In particular, if a quantity is useful, it is likely to have a name. The quantity [math]T = 1/2mv2, a scalar, is useful, so it gets a name: kinetic energy. The quantity 1/2mv2 , which is a vector whose magnitude is the kinetic energy and whose direction is the direction of motion, is not useful, so it doesn't have a name.

Now, there are at least three different reasons why kinetic energy, as a scalar, is a useful quantity in physics:
  1. It's part of the total energy---kinetic plus potential---which is very important because of conservation of energy. [math]E = T + V[/math] would have to be changed to [math]E = \|\mathbf{T}\| + V[/math] if we defined kinetic energy as a vector. That is, we'd just throw away the direction. There is no vector conservation law for energy, the way there is for momentum.
  2. It's a term in the Hamiltonian, which tells you how a physical system in general evolves in time. The Hamiltonian's value is usually the same as the total energy, [math]H = T + V[/math], although for some systems it's different. Again, if we defined kinetic energy as a vector, we'd just have to throw away the direction, [math]H = \|\mathbf{T}\| + V[/math].
  3. It's a term in the Lagrangian, which, like the Hamiltonian, governs the behaviour of a physical system, and is very important theoretically because its mathematical form makes certain properties of physical systems obvious. The Lagrangian for a particle is [math]L = T - V[/math]. Again, if we defined kinetic energy as a vector, we'd have to throw away the direction, and write [math]L = \|\mathbf{T}\| - V[/math].
There's no point in defining kinetic energy as a vector when we'd just have to throw away the direction every time we wanted to use it.
Akanksha Singh
7 years ago
Dear Dilip,
Can you clear your answer in simple and definite words. Its very difficult to understand this.
yaswanth r
7 years ago
workdone can be given by the expression F*S where F is the force acting and S is the displacement.Here,both F and S are vectors.We know the product of two vectors always gives us scalar.So,mathematically we can prove that Workdone is scalar.
yaswanth r
7 years ago
workdone can be given by the expression F*S where F is the force acting and S is the displacement.Here,both F and S are vectors.We know the product of two vectors always gives us scalar.So,mathematically we can prove that Workdone is scalar..
yaswanth r
7 years ago
workdone can be given by the expression F*S where F is the force acting and S is the displacement.Here,both F and S are vectors.We know the product of two vectors always gives us scalar.So,mathematically we can prove that Workdone is scalar...
maruthi sriharsha
6 years ago
energy=the work done by the body.
work=F.S                                     [since work is the dot product of force(F) and the displacement(S)]
as the dot product of two vectors gives a scalar but not a vector